OpenStudy (anonymous):

A research submarine has a 40.0 -diameter window 8.40 thick. The manufacturer says the window can withstand forces up to 1.20×106 . What is the submarine's maximum safe depth?

OpenStudy (anonymous):

The pressure inside the submarine is maintained at 1.0 atm

OpenStudy (anonymous):

Hi.Thanks

OpenStudy (anonymous):

Hi, i'd like to help, but can you specify the unit first?

OpenStudy (anonymous):

40 cm-8.40 cm-1.2 times 10^6 N

OpenStudy (anonymous):

I have no idea how to do this one,especially because it involves thickness

OpenStudy (anonymous):

Yes, the thickness doesn't matter. We need only force and area.

OpenStudy (anonymous):

Can youu explain more

OpenStudy (anonymous):

Well, it says 20 cm diameter window, what is window area? Then find maximum pressure=maximum force/window area.

OpenStudy (anonymous):

I got 0.50265 for area

OpenStudy (anonymous):

P=rho.g.h

OpenStudy (anonymous):

F=pghA

OpenStudy (anonymous):

1.2 time 10^6=.50265 times 1000 times 9.8 times h

OpenStudy (anonymous):

For h I got 243

OpenStudy (anonymous):

but it is wrong

OpenStudy (anonymous):

How did you find the area?

OpenStudy (anonymous):

Sorry I mean area is 012566

OpenStudy (anonymous):

but that didnt work out too

OpenStudy (anonymous):

I got 974.4

OpenStudy (anonymous):

you have to include pressure at sea level to find the pressure at any depth. p=p0+rho*g*h h=(p-p0)/rho*g where p is maximum pressure we talked before. You worked it this way?

OpenStudy (anonymous):

nope becuase the pressure inside is 1 atm

OpenStudy (anonymous):

I guess thickness might matter?

OpenStudy (anonymous):

I don't fin the relevance including thickness in this problem. It only matter if maximum force is not known. The thicker, the larger force it can handle.

OpenStudy (anonymous):

so what is wrong.Can you calculate and see what u get

OpenStudy (anonymous):

What is density of water used in this problem?

OpenStudy (anonymous):

1000 kg/m3 I guess

OpenStudy (anonymous):

or maybe I should use salt waters density?

OpenStudy (anonymous):

I got 946,21 for that density. If it's wrong too, try salt water.

OpenStudy (anonymous):

964,21

OpenStudy (anonymous):

946 is coorect

OpenStudy (anonymous):

not 964.what an accident ;ol

OpenStudy (anonymous):

You mean 964 is wrong?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

the answer is 947

OpenStudy (anonymous):

They used 1030 for seawater

OpenStudy (anonymous):

Try sea water, i think you'll get that number

OpenStudy (anonymous):

Thank you so much! I

OpenStudy (anonymous):

can you teach me how to solve pascal principle types of problems

OpenStudy (anonymous):

Heheh, you're welcome. I won't let my fan down :)

OpenStudy (anonymous):

like the hydrolic types

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):